# Sourced Carrollian Fluids Dual to Black Hole Horizons

### v2

@@ I Introduction
Fully unraveling the behavior of black holes in dynamical settings is key to our understanding of gravity. They lie at the core of some of the most extreme (potentially) observable phenomena in our universe and challenge our understanding on many fronts. Of particular current interest is to understand their nonlinear behavior at a fundamental level, even when it is tractable via numerical simulations. That is, to extract robust, first-principles information beyond specific cases.
@@
Nonlinear behavior of field theories is notoriously difficult to explore. Perturbative approaches in such regimes become unreliable or require high-order computations, which might render them impractical. Exploiting possible analogies with hydrodynamics—a field that has traditionally had to embrace the nonlinear regime through multiple means—provides us with complementary techniques to explore them. The early works by Damour and Znajek showed how black holes under electromagnetic fields can behave as ordinary conductors [1, 2] and how the components of the Einstein equations, when projected on the black hole horizon, can be interpreted as the equation of energy conservation and Navier–Stokes equations of a hypothetical horizon fluid [3]. These analogies were developed further within the membrane paradigm program [4, 5, 6, 7]. There, the fluid analogy was constructed through projections on a timelike surface (the stretched horizon) located outside the null event horizon. Such a timelike surface bypassed the singular nature of null surfaces and helped identify a fluid system with mechanical, electrical, and thermodynamic properties. The null nature of the event horizon obscured the structure of the resulting equations and was only recently interpreted as analogous to a special theory of fluids.
@@
A closely related link between hydrodynamics and gravity was pursued later with important results. Motivated by AdS/CFT holography [8], a duality between the long-wavelength perturbations of black holes in the bulk and relativistic hydrodynamics on the AdS boundary was derived [9, 10, 11, 12]. Subsequent works connected phenomena with gravitational dynamics, first in asymptotically AdS [13, 14, 15, 16] and later in asymptotically flat spacetimes [17, 18, 19]. Combined with the membrane paradigm, this duality has also been utilized for studying transport coefficients in boundary theory, drawing insights from gravity [20].
@@
In a recent parallel development, a duality of the properties of null surfaces and the vanishing speed of light, $c\rightarrow 0$, limit of relativistic hydrodynamics, named Carrollian fluids was shown [21]. This limit is the opposite of the Galilean limit, where the speed of light is taken to be infinite. In both cases, the relativistic description, which puts temporal and spatial coordinates on equal footing, no longer holds: the Galilean limit results in an absolute time, whereas the Carrollian limit yields an absolute space [22]. Remarkably, the resulting equations—upon a suitable identification—can be linked to those obtained by the projection of Einstein equations onto a null surface; that are, Raychaudhuri and Damour equations.
The study of the Carrollian limit began with the works of Lévy-Leblond [23] and Gupta [24] as a mathematical exploration of the contraction [25] of the Poincaré group, in which the speed of light vanishes. These studies evolved into an active field in physics in the last decade [26]. Some of the works on the field focused on the relations between the asymptotic symmetries of asymptotically flat spacetimes with Carroll groups and their use in flat holography [27, 28, 29, 30, 31]; Carrollian fluids as the limit of relativistic hydrodynamics [32, 33, 34, 35]; the interacting particle motion [36] and dynamics of extended objects [37]; Carrollian limits of relativistic field theories [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]  [^1] Note also the early work (1998) by Dautcourt [48] on the “ultrarelativistic” limit of general relativity.; applications in cosmology with emphasis on dark energy [49, 50]. Thermodynamic properties of Carrollian fluids have also been studied in recent works: for example, in [51] from the point of view of symmetries (as was developed in a general framework in [52]), and in [53] within a holographic setup by relating the gravitational radiation with the dissipation in the dual Carrollian theory on the boundary. Here, we focus on the special case of Carrollian fluids dual to black hole horizons and utilize our observations on the gravitational side of the duality to inform the Carrollian side.
It was later demonstrated that Damour’s preliminary ideas on the duality between the event horizon and Galilean fluids were naturally accommodated as a duality with Carrollian fluids [21, 54]. A dictionary between gravitational (geometrical) parameters and fluid parameters has been established [21, 55, 56]. The recent works by Jai-akson and Freidel [35, 57] generalized the structure of Carroll symmetries and developed a framework that is applicable for both null and timelike hypersurfaces. They used this formalism to treat the problems of the membrane paradigm by defining a stretched Carrollian structure on the timelike stretched horizons and obtaining the structure on the null horizon via taking a limit through the foliation of stretched horizons.
@@
Most of the work cited above on the Carrollian literature has so far focused on the kinematics of the vanishing-speed-of-light limit of relativistic theories and of null surfaces, in particular, the null infinity. While the notion of Carrollian fluids has also been studied as the limiting case of relativistic hydrodynamics, the Carrollian fluid dual to gravity has often been studied in vacuum spacetimes without a particular focus on the dynamics. Recently, Redondo-Yuste and Lehner [55] took a dynamical point of view and studied the Carrollian fluid dual of a spherically symmetric black hole perturbed by gravitational waves. They interpreted their results to connect the properties of the fluid, such as its viscosity, to the metric perturbations by adapting the dictionary between Carrollian fluid parameters and the geometric properties of the horizon.
@@
In this work, we study the Carrollian fluid/gravity duality in the non-vacuum case. We follow the stretched Carrollian geometry formalism laid out in [56, 57] with specific coordinate choices while incorporating matter fields in the bulk spacetime. After establishing the general form of the equations, we focus on a massless scalar field in the bulk as an instructive example. We then concentrate on an ideal system with a spherically symmetric black hole coupled to an infalling massless scalar field and solve the coupled equations for fluid parameters perturbatively to demonstrate the effects of the existence of the matter fields. Linking our results on the evolution of the horizon with the behavior of the dual Carrollian fluid, we connect the relaxation of the former to the equilibration of the latter.
@@
This work is organized as follows: In Section II, we lay out the geometric setup following the stretched Carrollian geometry formalism. We make judicial choices departing from the general formalism established in [57] to have a system of equations tractable with numerical methods. In Section III, we summarize the results of previous work on constructing the fluid dual to null surfaces, focusing specifically on the black hole horizon within our gauge choices. In Section IV, we focus on deploying sources in the bulk of the spacetime and establish how these sources enforce changes in the gravitational equations describing the null surface. In Section V, we carry out a computation focusing on a Schwarzschild black hole perturbed by a scalar field to illustrate the effects of sources. We do so by employing a perturbative scheme and solving the resulting equations numerically. We conclude with a summary of our results and discussions in Section VI.
Some details are included in the appendices. In Appendix A, we summarize the derivatives and connections relevant to the stretched Carrollian structure. In Appendix B, we expand on the gauge choices implemented in Section II. We leave some expressions in the derivations of perturbative equations to Appendix C, while expounding the numerical implementation in Appendix D. Finally, in Appendix E, we explicitly treat two well-known spacetimes in the structure introduced in Section II as concrete examples.
@@ II Geometry: Horizon Embedding
In this section, we follow the stretched Carrollian structure [56, 57] which makes use of the null rigging formalism [58]. The null rigging formalism allows a proper treatment of null surfaces embedded in an ambient spacetime while the stretched Carrollian structure generalizes the notion of Carrollian symmetries to timelike hypersurfaces. The combined formalism in [56, 57] allows us to construct a Carrollian structure induced from a bulk spacetime. [^2] To study the Carrollian structure on the null surfaces, one does not need to introduce the stretching in practice. For example, see [59, 60].
@@
Let $\mathscr{M}$ be a 4-dimensional spacetime manifold with signature (,+,+,+)(-,+,+,+), which admits a null hypersurface $\mathscr{N}$. [^3] The following treatment applies to Lorentzian manifolds of arbitrary dimensions. Here, we restrict ourselves to 4 dimensions for concreteness. Let $\mathscr{H}_{\rho}$ for $\rho>0$ denote any member of a foliation of 3-dimensional timelike hypersurfaces (or stretched horizons) embedded in $\mathscr{M}$ through an embedding function $r(x^{a})=\rho$, where $\rho$ is called the “stretching” [57]. Here $x^{a}$ denotes the coordinates on $\mathscr{M}$ and the letter $r$ is used for the foliation function to elucidate its radial characteristic. This foliation is constructed in a way that the null surface $\mathscr{N}$ is located at $r(x^{a})=0$. Thus, taking the limit $r(x^{a})=\rho\rightarrow 0$ corresponds to going from timelike hypersurfaces to the null surface in the “center” of the radial foliation i.e., $\mathscr{N}=\mathscr{H}_{\rho\rightarrow 0}$. In the following, we will use $\mathscr{H}$ dropping the subscript $\rho$ to denote a leaf of the foliation for which $r(x^{a})=\rho$.
@@
A weak Carroll structure can be defined as a fiber bundle directly on the null hypersurface $\mathscr{N}$ if a null metric of the base manifold, $\mathcal{S}$, and an appropriate vector field is introduced on it [61, 29, 35]. When this vector field is interpreted as a time coordinate, the fiber bundle structure represents the splitting of time and space coordinates. [^4] See also [22] for the relation between Carrollian structures to Newton-Cartan geometries. On the timelike hypersurface $\mathscr{H}$, the stretched Carrollian structure can be defined by appropriate choices of a Carrollian vector field, Ehresmann connection, the induced metric, and the stretching $\rho$. [^5] Further details on the construction of Carroll structures can be found in the literature [35, 29]. Next, we introduce these geometric ingredients for the Carrollian structure.
@@ II.1 Vectors, Forms & Projectors
Here, we summarize the recipe for the stretched Carrollian structure presented in [57] for completeness. We then restrict ourselves to a particular gauge choice in the subsequent sections.
In the following, we reserve the lowercase Latin indices from the first half of the alphabet to denote the indices of tensors defined on $\mathscr{M}$. We use the letters starting from $i$ for the indices intrinsic to the coordinates on $\mathscr{H}$. Finally, we use capital Latin letters to denote the coordinates intrinsic to the sphere manifold $\mathcal{S}$, which is the spacelike base manifold of the null hypersurface $\mathscr{N}$. Unless otherwise specified, we work with geometric units in which $c=G=1$.
We denote the metric on $\mathscr{M}$ by $g_{ab}$ and the covariant derivative compatible with it by $\nabla_{a}$. We also define on $\mathscr{M}$: a vector field $l^{a}$ which is tangent to $\mathscr{H}$, a normal form $n_{a}$, and a null rigging vector $k^{a}$. These objects, alongside their metric duals, satisfy the following relations:
@@
Restricting ourselves to the indices intrinsic to $\mathscr{H}$, we now have established the stretched Carrollian structure on $\mathscr{H}$ with the data $\left(l^{i},k_{i},h_{ij},\rho\right)$ [57, 56].
The stretched Carrollian structure on the timelike hypersurface connects naturally to the Carrollian structure on the null hypersurface for the vanishing stretching $\rho$. To see this, we can decompose the metric on $\mathscr{H}$ as the following
@@
This projector further projects the tensors on $\mathscr{M}$ to the horizontal base space of the bundle, $\mathcal{S}$. Using the horizontal projector, we can also induce the null Carrollian metric directly from the metric in the bulk as
@@
The hypersurface $\mathscr{N}$ has the structure of a fiber bundle with a 2-dimensional base $\mathcal{S}$ and a one-dimensional vector field $l^{a}$ as its fiber [29]. In this respect, the 1-form $k_{a}$ is also referred to as the Ehresmann connection, which allows a splitting of vertical (along $l^{a}$) and horizontal directions in the fiber bundle [56]. Through $k_{a}$, a basis frame $\left(e_{A}\right)^{a}$ is defined on the horizontal space, which satisfies $\left(e_{A}\right)^{a}k_{a}=0$. The basis of the horizontal subspace of $\mathscr{N}$ can be expressed with the frame $\left(e_{A}\right)^{a}$ and co-frame field $\left(e_{A}\right)_{a}$ satisfying $q_{AB}=\left(e_{A}\right)_{a}\left(e_{B}\right)^{a}$. Here, the uppercase Latin indices are frame indices and are raised and lowered via $q_{AB}$, which is the metric on the sphere $\mathcal{S}$.
@@
where the square brackets denote the anti-symmetrization of indices. The first quantity, $\varphi_{i}$, is the Carrollian acceleration; whereas, the second one, $\varpi_{ij}$, is the Carrollian vorticity [21]. These quantities also appear in the commutation relations of the vectors defined above, as demonstrated in [56].
@@
This quantity is interpreted as an “surface density of impulsion” [62], or fluid momentum by Damour [3]. Later, we will demonstrate the relation between the Hájíček one-form and the horizon angular momentum in a dynamical setting.
The deviation tensors with their components, surface gravity, and the Hájíček field are the essential pieces for the fluid description of the horizon. As we will discuss in the upcoming sections, the stress-energy tensor describing $\mathscr{N}$, which admits the Carrollian fluid interpretation, is formulated using these geometric quantities. Moreover, the equations describing the dynamics of the horizon surface are conveniently described in terms of these quantities.

### v3

@@ I Introduction
Fully unraveling the behavior of black holes in dynamical settings is key to our understanding of gravity. They lie at the core of some of the most extreme (potentially) observable phenomena in our universe and challenge our understanding on many fronts. Of particular current interest is understanding their nonlinear behavior at a fundamental level, even when it is tractable via numerical simulations. That is, to extract robust, first-principles information beyond specific cases.
The nonlinear behavior of field theories is notoriously difficult to explore. Perturbative approaches in such regimes become unreliable or require high-order computations, which might render them impractical. Exploiting possible analogies with hydrodynamics, a field that has traditionally had to embrace the nonlinear regime through multiple means, provides us with complementary techniques to explore them. In this work, we take this route to further this connection in gravity while also shedding light on a particular set of hydrodynamical equations.
@@
The early work by Damour showed how the components of the Einstein equations, when projected on the black hole horizon, can be interpreted as the equation of energy conservation and the Navier–Stokes equations of a hypothetical horizon fluid [Damour1982]. These analogies were further developed within the membrane paradigm program [TPM1986, Price1986, Suen1985, Thorne1982]. There, the fluid analogy was constructed through projections on a timelike surface (the stretched horizon) located outside the null event horizon. Such a timelike surface bypassed the singular nature of null surfaces and helped identify a fluid system with mechanical, electrical, and thermodynamic properties. The null nature of the event horizon obscured the structure of the resulting equations and was only recently interpreted as analogous to a special theory of fluids.
@@
A closely related link between hydrodynamics and gravity was later pursued with important results. Motivated by AdS/CFT holography [Maldacena_1999], a duality was derived between long-wavelength perturbations of black holes in the bulk and relativistic hydrodynamics on the AdS boundary [Policastro_2002, Kovtun2003, Bhattacharyya_2008, hubeny2011fluidgravitycorrespondence]. Subsequent work connected phenomena with gravitational dynamics, first in asymptotically AdS [Carrasco_2012, Adams_2014, Green_2014, Raamsdonk_2008] and later in asymptotically flat spacetimes [Yang_2015, Iuliano:2024ogr, Ma:2025rnv]. Combined with the membrane paradigm, this duality has also been utilized to study transport coefficients in boundary theory, drawing insights from gravity [Iqbal_2009].
@@
In a more recent development, a duality of the properties of null surfaces and the vanishing speed of light, $c\rightarrow 0$ limit of relativistic hydrodynamics, named Carrollian fluids was shown [Donnay_2019]. This limit is opposite to the Galilean limit, where the speed of light is taken to be infinite. In both cases, the relativistic description, which puts temporal and spatial coordinates on equal footing, no longer holds: the Galilean limit results in an absolute time, whereas the Carrollian limit yields an absolute space [Duval_2014]. Remarkably, the resulting equations—upon a suitable identification—can be linked to those obtained by the projection of Einstein equations onto a null surface; that are, Raychaudhuri and Damour equations.
@@
Most of the work on the Carrollian fluid–gravity connection has so far focused on the kinematics of the vanishing-speed-of-light limit of relativistic theories and of null surfaces, in particular, the null infinity. While the notion of Carrollian fluids has also been studied as the limiting case of relativistic hydrodynamics, the Carrollian fluid dual to gravity has often been studied in vacuum spacetimes without a particular focus on the dynamics. Recently, Redondo-Yuste and Lehner [Redondo_Yuste_2023] took a dynamical point of view and studied the Carrollian fluid dual of a spherically symmetric black hole perturbed by gravitational waves. They interpreted their results to connect the properties of the fluid, such as its viscosity, to the metric perturbations by adapting the dictionary between Carrollian fluid parameters and the geometric properties of the horizon.
@@
In this work, we study the Carrollian fluid/gravity duality in the nonvacuum case. We follow the stretched Carrollian geometry formalism laid out in [freidel2022carrollian, freidel2024geometrycarrollianstretchedhorizons] with specific coordinate choices while incorporating matter fields in the bulk spacetime. After establishing the general form of the equations, we focus on a massless scalar field in the bulk as an instructive example. We then concentrate on an ideal system with a spherically symmetric black hole coupled to an infalling massless scalar field and solve the coupled equations for fluid parameters perturbatively to demonstrate the effects of the existence of the matter fields. We carry out perturbative computations up to second order in the scalar field strength as it is the lowest order at which the matter terms affect the horizon geometry. Linking our results on the evolution of the horizon with the behavior of the dual Carrollian fluid, we connect the relaxation of the former to the equilibration of the latter.
@@
This work is organized as follows: In Section II, we lay out the geometric setup following the stretched Carrollian geometry formalism. We make judicial choices departing from the general formalism established in [freidel2024geometrycarrollianstretchedhorizons] to have a system of equations that is tractable with numerical methods. In Section III, we summarize the results of previous work on constructing fluid dual to null surfaces, focusing specifically on the black hole horizon within our gauge choices. In Section IV, we focus on the deployment of sources in the bulk of spacetime and establish how these sources enforce changes in the gravitational equations describing the null surface. In Section V, we carry out a computation focusing on a Schwarzschild black hole perturbed by a scalar field to illustrate the effects of sources. We do so by employing a perturbative scheme and solving the resulting equations numerically. We conclude with a summary of our results and discussions in Section VI.
@@ II Geometry: Horizon Embedding
The duality between the event horizon and fluids were naturally accommodated as a duality with Carrollian fluids [Donnay_2019, Adami_2021]. A dictionary has been established between gravitational (geometrical) parameters and fluid parameters [Donnay_2019, Redondo_Yuste_2023, freidel2022carrollian]. The recent work by Jai-akson and Freidel [freidel2022carrollian, Freidel_2023, freidel2024geometrycarrollianstretchedhorizons] generalized the structure of Carroll symmetries and developed a framework that is applicable for both null and timelike hypersurfaces. They used this formalism to treat the problems of the membrane paradigm by defining a stretched Carrollian structure on the timelike stretched horizons and obtaining the structure on the null horizon via taking a limit through the foliation of stretched horizons.
@@
In this section, we follow the stretched Carrollian structure [freidel2022carrollian, freidel2024geometrycarrollianstretchedhorizons] which makes use of the null rigging formalism [Mars_1993]. The null rigging formalism allows a proper treatment of null surfaces embedded in an ambient spacetime, while the stretched Carrollian structure generalizes the notion of Carrollian symmetries to timelike hypersurfaces. The combined formalism in [freidel2022carrollian, freidel2024geometrycarrollianstretchedhorizons] allows us to construct a Carrollian structure induced from a bulk spacetime. [^1] To study the Carrollian structure on the null surfaces, one does not need to introduce the stretching in practice. For example, see [Ciambelli:2023mvj, ciambelli2023nullraychaudhuricanonicalstructure].
@@
Let $\mathscr{M}$ be a 4-dimensional spacetime manifold with signature (,+,+,+)(-,+,+,+), which admits a null hypersurface $\mathscr{N}$. [^2] The following treatment applies to Lorentzian manifolds of arbitrary dimensions. Here, we restrict ourselves to 4 dimensions for concreteness. Let $\mathscr{H}_{\rho}$ for $\rho>0$ denote any member of a foliation of 3-dimensional timelike hypersurfaces (or stretched horizons) embedded in $\mathscr{M}$ through an embedding function $r(x^{a})=\rho$, where $\rho$ is called the “stretching” [freidel2024geometrycarrollianstretchedhorizons]. Here, $x^{a}$ denotes the coordinates on $\mathscr{M}$ and the letter $r$ is used for the foliation function to elucidate its radial characteristic. This foliation is constructed so that the null surface $\mathscr{N}$ is located at $r(x^{a})=0$. Thus, taking the limit $r(x^{a})=\rho\rightarrow 0$ corresponds to going from timelike hypersurfaces to the null surface in the “center” of the radial foliation, i.e., $\mathscr{N}=\mathscr{H}_{\rho\rightarrow 0}$. In the following, we will use $\mathscr{H}$ dropping the subscript $\rho$ to denote a leaf of the foliation for which $r(x^{a})=\rho$.
@@
A weak Carroll structure can be defined as a fiber bundle directly on the null hypersurface $\mathscr{N}$ if a null metric of the base manifold, $\mathcal{S}$, and an appropriate vector field is introduced on it [Duval_2014_3, Ciambelli_2019, Freidel_2023]. When this vector field is interpreted as a time coordinate, the fiber bundle structure represents the splitting of time and space coordinates. [^3] See also [Duval_2014] for the relation between Carrollian structures to Newton-Cartan geometries. On the timelike hypersurface $\mathscr{H}$, the stretched Carrollian structure can be defined by appropriate choices of a Carrollian vector field, the Ehresmann connection, the induced metric, and the stretching $\rho$. [^4] Further details on the construction of Carroll structures can be found in the literature [Freidel_2023, Ciambelli_2019]. Next, we introduce these geometric ingredients for the Carrollian structure.
@@ II.1 Vectors, Forms & Projectors
Here, we summarize the recipe for the stretched Carrollian structure presented in [freidel2024geometrycarrollianstretchedhorizons] for completeness. We then restrict ourselves to a particular gauge choice in the subsequent sections.
In the following, we reserve the lowercase Latin indices from the first half of the alphabet to denote the indices of tensors defined on $\mathscr{M}$. We use letters starting from $i$ for the intrinsic indices to the coordinates in $\mathscr{H}$. Finally, we use capital Latin letters to denote the coordinates intrinsic to the sphere manifold $\mathcal{S}$, which is the spacelike base manifold of the null hypersurface $\mathscr{N}$. Unless otherwise specified, we work with geometric units in which $c=G=1$.
We denote the metric on $\mathscr{M}$ by $g_{ab}$ and the covariant derivative compatible with it by $\nabla_{a}$. We also define on $\mathscr{M}$: a vector field $l^{a}$ that is tangent to $\mathscr{H}$, a normal form $n_{a}$, and a null rigging vector $k^{a}$. These objects, along with their metric duals, satisfy the following relations:
@@
Restricting ourselves to the indices intrinsic to $\mathscr{H}$, we now have established the stretched Carrollian structure on $\mathscr{H}$ with the data $\left(l^{i},k_{i},h_{ij},\rho\right)$ [freidel2024geometrycarrollianstretchedhorizons, freidel2022carrollian].
The stretched Carrollian structure on the timelike hypersurface connects naturally to the Carrollian structure on the null hypersurface for the vanishing stretching $\rho$. To see this, we can decompose the metric on $\mathscr{H}$ as the following.
@@
This projector also projects the tensors on $\mathscr{M}$ to the horizontal base space of the bundle, $\mathcal{S}$. Using the horizontal projector, we can also induce the null Carrollian metric directly from the metric in the bulk, as
@@
The hypersurface $\mathscr{N}$ has the structure of a fiber bundle with a 2-dimensional base $\mathcal{S}$ and a one-dimensional vector field $l^{a}$ as its fiber [Ciambelli_2019]. In this respect, the 1-form $k_{a}$ is also known as the Ehresmann connection, which allows the separation of vertical (along $l^{a}$) and horizontal directions in the fiber bundle [freidel2022carrollian]. Through $k_{a}$, a basis frame $\left(e_{A}\right)^{a}$ is defined on the horizontal space, which satisfies $\left(e_{A}\right)^{a}k_{a}=0$. The basis of the horizontal subspace of $\mathscr{N}$ can be expressed with the frame $\left(e_{A}\right)^{a}$ and co-frame field $\left(e_{A}\right)_{a}$ satisfying $q_{AB}=\left(e_{A}\right)_{a}\left(e_{B}\right)^{a}$. Here, the uppercase Latin indices are frame indices and are raised and lowered through $q_{AB}$, which is the metric on the sphere $\mathcal{S}$.
@@
where the square brackets denote the anti-symmetrization of indices. The first quantity, $\varphi_{i}$, is the Carrollian acceleration; while the second quantity, $\varpi_{ij}$, is the Carrollian vorticity [Donnay_2019]. These quantities also appear in the commutation relations of the vectors defined above, as demonstrated in [freidel2022carrollian].
@@
This quantity is interpreted as an “surface density of impulsion” [Damour:1979wya], or fluid momentum by Damour [Damour1982]. Later, we will demonstrate the relation between the Hájíček one-form and the horizon angular momentum in a dynamical setting.
The deviation tensors with their components, surface gravity, and the Hájíček field are the essential pieces for the fluid description of the horizon. As we shall discuss in the next sections, the stress-energy tensor describing $\mathscr{N}$, which admits the Carrollian fluid interpretation, is formulated using these geometric quantities. Moreover, the equations describing the dynamics of the horizon surface are conveniently described in terms of these quantities.